A gambler rolls two dice and stands to loss Rs.2/- if he fails to throw a six ; to win
Rs.4/- he must throw one six and to win Rs.10/- he must throw 2 sixes. Is the game
fair?

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

Total number of possibilities on two dice
36
Number of possibilities with 6 on both dice
1
Number of possibilities with 6 on one dice
10
Number of possibilities with 6 on neither of dice
25
So lets calculate the amount of the money
25\times 2 = Rs.50
(10\times 4)+(1\times 10) = Rs.50
Both amount are same for winning and losing.
So the game is fair.